# Process Reinforcement through Implicit Rewards

### v1

@@ Abstract
Dense process rewards have proven a more effective alternative to the sparse outcome-level rewards in the inference-time scaling of large language models (LLMs), particularly in tasks requiring complex multi-step reasoning. While dense rewards also offer an appealing choice for the reinforcement learning (RL) of LLMs since their fine-grained rewards have the potential to address some inherent issues of outcome rewards, such as training efficiency and credit assignment, this potential remains largely unrealized. This can be primarily attributed to the challenges of training process reward models (PRMs) online, where collecting high-quality process labels is prohibitively expensive, making them particularly vulnerable to reward hacking. To address these challenges, we propose PRIME (Process Reinforcement through IMplicit rEwards), which enables online PRM updates using only policy rollouts and outcome labels through implict process rewards. PRIME combines well with various advantage functions and forgoes the dedicated reward model training phase that existing approaches require, substantially reducing the development overhead. We demonstrate PRIME’s effectiveness on competitional math and coding. Starting from Qwen2.5-Math-7B-Base, PRIME achieves a 15.1% average improvement across several key reasoning benchmarks over the SFT model. Notably, our resulting model, Eurus-2-7B-PRIME, surpasses Qwen2.5-Math-7B-Instruct on seven reasoning benchmarks with 10% of its training data. [^1] Models and data are available at: https://github.com/PRIME-RL/PRIME.
@@ 1 Introduction
Dense process rewards, which provide feedback at each intermediate step rather than only the whole trajectory, have proven effective in inference-time scaling of large language models (LLMs) on challenging reasoning tasks (Uesato et al., 2022; Lightman et al., 2023; Wang et al., 2023; Yuan et al., 2024b). On the training side, they also present superiorities in the reinforcement learning (RL) of LLMs, particularly in improving training efficiency (Sutton & Barto, 2018) and credit assignment (Leike et al., 2018) compared with sparse outcome rewards. However, successful applications of dense rewards in RL for LLMs are limited (Setlur et al., 2024), as current industry-leading models primarily depend on verifiable outcome rewards and have not yet demonstrated meaningful progress with dense rewards (DeepSeek-AI et al., 2025; Team et al., 2025).
We identify the central challenge as how to acquire and utilize high-quality dense rewards at scale, which enables online process reward model (PRM) update efficiently. The reason is that, optimizing towards a static reward model eventually leads to overoptimization or reward hacking (Gao et al., 2022) due to distribution shift. Ideally, this can be solved by improving the reward model online (Leike et al., 2018). However, acquiring dense process labels for training is prohibitively more expensive. Existing methods either need to build complicated human annotation pipelines (Lightman et al., 2023) or rely on estimation-based methods, which require about 10$\times$ more rollouts for each step than sampling only the response-level trajectories (Wang et al., 2023; Kazemnejad et al., 2024). Neither of them is scalable in online RL. Moreover, to the best of our knowledge, it remains underexplored how to incorporate dense rewards into RL for LLMs.
@@
In this work, we propose Process Reinforcement through Implicit Rewards (PRIME), a scalable framework for enhancing reasoning capabilities via efficient reinforcement learning with dense token-level rewards. At its core, the framework employs recently proposed implicit process reward modeling (Yuan et al., 2024b) to train dense reward models with only outcome-level labels. This enables PRIME to perform online learning of reward signals using only outcome labels on policy rollouts, thereby fundamentally mitigating reward hacking while maintaining the same computational cost as traditional outcome reward models (ORMs). Besides scalability, PRIME also (1) serves as a general method to fuse token-level dense rewards and sparse outcome rewards by calculating their returns separately before summing together, which is compatible with diverse RL algorithms (Williams, 1992; Kool et al., 2019; Shao et al., 2024; Ahmadian et al., 2024; Schulman et al., 2017); (2) eliminates the dedicated reward modeling stage, which is required by existing works, by simply initializing from the SFT model or even the base model (§ 5.6). In summary, starting from one single language model, the PRIME framework can efficiently accomplish the generation of dense rewards, the initialization and updating of reward models, as well as the reinforcement learning (RL) training of the policy model.
Model Eurus-2-7B-PRIME Qwen2.5-Math-7B-Instruct
Base Model Qwen2.5-Math-7B Qwen2.5-Math-7B
SFT Data 230K (open-source) 2.5M (open-source & in-house)
RM Data 0 618K (in-house)
RM Eurus-2-7B-SFT Qwen2.5-Math-RM (72B)
RL Data 150K queries $\times$ 4 samples 66K queries $\times$ 32 samples
@@
In experiments, we train Qwen2.5-Math-7B-Base (Yang et al., 2024b) with PRIME after a lightweight SFT warmup stage. Compared to RL using outcome rewards only, PRIME achieves a 2.5×2.5$\times$2.5 × sample efficiency gain and a $6.9\%$ performance improvements on challenging math problems. As shown in Figure 1, through PRIME, we successfully achieve substantial improvement on key mathematical reasoning benchmarks over the SFT model, leading to 16.7% improvement on average, and over 20% on AMC&AIME competitions. Our final model Eurus-2-7B-PRIME surpassed Qwen2.5-Math-7B-Instruct on five key mathematical benchmarks. Notably, this is achieved with only $10\%$ of the data used by Qwen-Math, as in Table 1.
@@
Our analysis shows that updating the PRM online is key to the success of PRIME (§5.1). We also show that PRIME could generally boost various RL algorithms, including RLOO (Ahmadian et al., 2024), REINFORCE (Williams, 1992), PPO (Schulman et al., 2017), and GRPO (Shao et al., 2024)5.4). In terms of the design choices of advantage estimate, we observe that Implicit PRMs are better to be used as reward models than value models (§5.5).
@@ 2.1 RL Preliminaries for LLMs
$\gamma\in[0,1]$ is a discount factor that optionally decays future rewards, and $r(y_{s})$ is the reward provided by the environment at time step $s$ with $x$ and $\mathbf{y}_{<s}$ being omitted in conditions. Eq. 2 is the general formula of the Monte-Carlo (MC) advantage estimate, which indicates that, the high-quality and dense reward at each step is crucial for RL. Different choices of $b$ include, e.g. directly using values Williams (1992), group average of rewards (Shao et al., 2024), and leave-one-out average of rewards Ahmadian et al. (2024); Kool et al. (2019).
@@ 2.2 Key Challenges in Scalable Dense Rewards
Notably, a concurrent work shares similar conclusions and thus is impeded from incorporating PRMs into their large-scale RL training (DeepSeek-AI et al., 2025).
@@ 3.1 Enabling Scalable Reward Update with Implicit Reward Modeling
Input Language model $\pi_{\theta_{\text{init}}}$; outcome reward verifier $r_{o}$; dataset $\mathcal{D}$; sample number $K$; total iteration $N$.
1:Initialize policy model $\pi_{\theta}\leftarrow\pi_{\theta_{\text{init}}}$, $\pi_{\theta_{\text{old}}}\leftarrow\pi_{\theta_{\text{init}}}$, implicit PRM $\pi_{\phi}\leftarrow\pi_{\theta_{\text{init}}}$, reference model $\pi_{\text{ref}}\leftarrow\pi_{\theta_{\text{init}}}$
@@
9:     Compute advantages $A$ with Eq. 5
@@
In PRIME, upon rollouts being generated and graded by the (ground truth) outcome verifier, we update the Implicit PRM online with on-policy rollouts and outcome supervision and then calculate token-level dense rewards to estimate advantages, which solves C1 and C2 mentioned in §2.2 respectively: (1) To prevent overoptimization and reward hacking, it is crucial to update reward models online. However, updating previous PRMs (Lightman et al., 2023) requires annotating step labels on the latest policy rollouts, which is neither efficient nor scalable during online RL. In contrast, the Implicit PRM only demands outcome labels to train due to its special reward representation, and thus it can be easily updated with policy rollouts and outcome labels or rewards, both of which have already been collected to update the policy model. (2) Unlike common PRMs that produce only step-level rewards, the Implicit PRM provides more fine-grained token-level rewards at no additional cost. This addresses the ambiguity in identifying steps in LLM responses while not introducing extra overhead, making it easy to combine with any RL algorithms for advantage estimation.
@@ 3.2 Advantage Estimation and Policy Update
$A^{i}=r(\mathbf{y}^{i}_{T})-\frac{1}{K-1}\sum_{j\neq i}r(\mathbf{y}^{j}_{T})$ (4)
where $r(\mathbf{y}^{i}_{T})$ denotes the reward of $i$-th response at final step $T$, $K$ is the number of samples for one prompt. The leave-one-out (LOO) baseline helps reduce variances.
@@
Updating policy with PPO clip surrogate loss. We adopt PPO clip surrogate loss for more stable policy updates:
@@
where $\epsilon$ is a clipping parameter. The loss prevents the updated policy from deviating too far from the original distribution, which is the prerequisite of importance sampling. The legitimacy of importance sampling then enables the reuse of rollouts sampled in previous steps, thus improving sampling efficiency.
@@ 3.3 Other Techniques
Initializing PRM with SFT/base model. In practice, we find that the starting policy model itself serves as a decent initialization of PRM, bypassing the PRM training stage. This solves C3 in §2.2 and even outperforms a dedicatedly trained PRM, as shown in § 5.1.
Figure 3: Impact of online prompt filtering on training rewards.
@@ 4 Experiments
4.1 Imitation Warmup
We focus on mathematical and coding problems in this paper. For models, we start with Qwen2.5-Math-7B-Base (Yang et al., 2024b) for its great mathematical capabilities. We first performed supervised finetuning for RL preparation.
Data Construction. To construct the SFT dataset, we collect reasoning instructions from several open-source datasets. For completion, we employed LLaMA-3.1-70B-Instruct (Meta, 2024) to answer the instructions, with a system prompt requesting the model to perform action-centric chain-of-thought. We finally obtained 230K SFT data, the detailed sources and statistics can be found in § A.
@@ 4.2 RL Settings
Rule-based Outcome Verifier. Consistent with recent research that adopts exact match with ground truth as unhackable rewards (Gao et al., 2024; Lambert et al., 2024; DeepSeek-AI et al., 2025), we define the rule-based ground truth outcome verifiers (OV) for math and coding as follows:
$r_{o}^{\text{math}}(\mathbf{y})$ =$=\begin{cases}1,&\text{matched}\\ 0,&\text{otherwise}\end{cases}\quad r_{o}^{\text{code}}(\mathbf{y})=\frac{\sum\text{\#passes}}{\sum\text{\#test cases}}$
@@
Table 2: Detailed results of PRIME and RLOO w/ outcome verifier (OV). At the same 240 steps, the model trained by PRIME is generally better than the model trained by outcome rewards.
@@
Figure 4: The effect of dense reward. We compare PRIME and RLOO with outcome verifier (OV). Dense rewards in PRIME lead to 2.5×2.5$\times$2.5 × sample efficiency and $6.9\%$ performance improvement. PRIME also substantially outperforms RLOO on downstream tasks.
@@ 4.3 Main Results
As shown in Figure 1 and Table 2, Eurus-2-7B-PRIME achieves substantial improvements on key reasoning benchmarks over the SFT version of the model, leading to 15.1% improvement on average, and over 20% on AMC and AIME competitions. Besides, Eurus-2-7B-PRIME achieves 26.7% pass@1 on AIME 2024, surpassing GPT-4o, Llama-3.1-70B-Instruct, and Qwen2.5-Math-7B-Instruct, demonstrating its excellent reasoning ability.
@@
We first validate the effect of dense rewards compared to RLOO with outcome rewards only. We train this model for 240 steps. For PRIME, we use the same setting and train the model for 592 steps. We plot the training rewards measured by the outcome verifier and test accuracy in Figure 4. Compared with sparse reward, PRIME takes 40% of the training steps to achieve the same training rewards as RLOO and improves the final rewards by 6.9%, with lower variances. On downstream tasks, PRIME also consistently outperforms OV only setup. Detailed results are listed in Table 2.
@@ 5 Analysis
Figure 5: Comparison of different PRMs. Online PRM initialized from SFT model achieved the best results. Surprisingly, using PRMs trained on extra rollouts hurts the performance in both online and offline settings.
@@ 5.1 Design Choices for the Implicit PRM
SFT model initializes a good PRM. Conventionally, we need to collect data to train RMs and PRMs, and then we can use them in RL. However, the Implicit PRM is a language model, so we can initialize it from any language model with the same tokenizer as the policy model. To investigate whether it is still necessary to train a PRM in advance, we conduct experiments with different PRM initialization strategies: with the SFT model itself and with a specially trained PRM. For the later one, we train EurusPRM from Eurus-2-7B-SFT with additional 500K data generated by Llama3.1 and Qwen2.5 series (data details in § B.5).
Figure 6: Impact of PRM online update. The offline PRM is gradully been overoptimized while online PRMs achieve higher accuracy throughout training.
@@
5.2 Reference Model Choice is Flexible

### v2

@@ Abstract
Dense process rewards have proven a more effective alternative to the sparse outcome-level rewards in the inference-time scaling of large language models (LLMs), particularly in tasks requiring complex multi-step reasoning. While dense rewards also offer an appealing choice for the reinforcement learning (RL) of LLMs since their fine-grained rewards have the potential to address some inherent issues of outcome rewards, such as training efficiency and credit assignment, this potential remains largely unrealized. This can be primarily attributed to the challenges of training process reward models (PRMs) online, where collecting high-quality process labels is prohibitively expensive, making them particularly vulnerable to reward hacking. To address these challenges, we propose PRIME (Process Reinforcement through IMplicit rEwards), which enables online PRM updates using only policy rollouts and outcome labels through implicit process rewards. PRIME combines well with various advantage functions and forgoes the dedicated reward model training phase that existing approaches require, substantially reducing the development overhead. We demonstrate PRIME’s effectiveness on competitional math and coding. Starting from Qwen2.5-Math-7B-Base, PRIME achieves a 15.1% average improvement across several key reasoning benchmarks over the SFT model. Notably, our resulting model, Eurus-2-7B-PRIME, surpasses Qwen2.5-Math-7B-Instruct on seven reasoning benchmarks with 10% of its training data.
@@ 1 Introduction
Dense process rewards, which provide feedback at each intermediate step rather than only the whole trajectory, have proven effective in inference-time scaling of large language models (LLMs) on challenging reasoning tasks (Uesato et al., 2022; Lightman et al., 2023; Wang et al., 2023; Yuan et al., 2024b). On the training side, they also present superiority in the reinforcement learning (RL) of LLMs, particularly in improving training efficiency (Sutton & Barto, 2018) and credit assignment (Leike et al., 2018) compared with sparse outcome rewards. However, successful applications of dense rewards in RL for LLMs are limited (Setlur et al., 2024), as current industry-leading models primarily depend on verifiable outcome rewards and have not yet demonstrated meaningful progress with dense rewards (DeepSeek-AI et al., 2025; Team et al., 2025).
[…]
@@
In this work, we propose Process Reinforcement through Implicit Rewards (PRIME), a scalable framework for enhancing reasoning capabilities via efficient reinforcement learning with dense token-level rewards. At its core, the framework employs recently proposed implicit process reward modeling (Yuan et al., 2024b) to train dense reward models with only outcome-level labels. This enables PRIME to perform online learning of reward signals using only outcome labels on policy rollouts, thereby fundamentally mitigating reward hacking while maintaining the same computational cost as traditional outcome reward models (ORMs). Besides scalability, PRIME also (1) serves as a general method to fuse token-level dense rewards and sparse outcome rewards by calculating their returns separately before summing together, which is compatible with diverse RL algorithms (Williams, 1992; Kool et al., 2019; Shao et al., 2024; Ahmadian et al., 2024; Schulman et al., 2017); (2) eliminates the dedicated reward modeling stage, which is required by existing works, by simply initializing from the SFT model or even the base model (§ B.4). In summary, starting from one single language model, the PRIME framework can efficiently accomplish the generation of dense rewards, the initialization and updating of reward models, as well as the reinforcement learning (RL) training of the policy model.
@@
In experiments, we train Qwen2.5-Math-7B-Base (Yang et al., 2024b) with PRIME after a lightweight SFT warmup stage. Compared to RL using outcome rewards only, PRIME achieves a $2.5\times$ sample efficiency gain and a $6.9\%$ performance improvements on challenging math problems. As shown in Figure 12, through PRIME, we successfully achieve substantial improvement on key mathematical reasoning benchmarks over the SFT model, leading to 16.7% improvement on average, and over 20% on AMC&AIME competitions. Our final model Eurus-2-7B-PRIME surpassed Qwen2.5-Math-7B-Instruct on five key mathematical benchmarks. Notably, this is achieved with only $10\%$ of the data used by Qwen-Math, as in Table 3.
@@
Our analysis shows that updating the PRM online is key to the success of PRIME (§5.1). We also show that PRIME could generally boost various RL algorithms, including RLOO (Ahmadian et al., 2024), REINFORCE (Williams, 1992), PPO (Schulman et al., 2017), and GRPO (Shao et al., 2024)5.3). In terms of the design choices of advantage estimate, we observe that Implicit PRMs are better to be used as reward models than value models (§5.4).
@@ 2.1 RL Preliminaries for LLMs
$\gamma\in[0,1]$ is a discount factor that optionally decays future rewards, and $r(y_{s})$ is the reward provided by the environment at time step $s$ with $x$ and $\mathbf{y}_{<s}$ being omitted in conditions. Eq. 2 is the general formula of the Monte-Carlo (MC) advantage estimate, which indicates that, the high-quality and dense reward at each step is crucial for RL. Different choices of $b$ include, e.g. directly using values (Williams, 1992), group average of rewards (Shao et al., 2024), and leave-one-out average of rewards (Ahmadian et al., 2024; Kool et al., 2019).
@@ 2.2 Key Challenges in Scalable Dense Rewards
Notably, DeepSeek-AI et al. (2025) shares similar conclusions and thus is impeded from incorporating PRMs into large-scale RL training.
@@ 3.1 Enabling Scalable Reward Update with Implicit Reward Modeling
In PRIME, upon rollouts being generated and graded by the (ground truth) outcome verifier, we update the Implicit PRM online with on-policy rollouts and outcome supervision and then calculate token-level dense rewards to estimate advantages, which solves C1 and C2 mentioned in §2.2 respectively: (1) To prevent overoptimization and reward hacking, it is crucial to update reward models online. However, updating previous PRMs (Lightman et al., 2023) requires annotating step labels on the latest policy rollouts, which is neither efficient nor scalable during online RL. In contrast, the Implicit PRM only demands outcome labels to train due to its special reward representation, and thus it can be easily updated with policy rollouts and outcome labels or rewards, both of which have already been collected to update the policy model.
@@
Input Language model $\pi_{\theta_{\text{init}}}$; outcome verifier $r_{o}$; dataset $\mathcal{D}$; sample number $K$; total iteration $N$.
1:Initialize policy model $\pi_{\theta},\pi_{\theta_{\text{old}}}\leftarrow\pi_{\theta_{\text{init}}}$, implicit PRM and reference model $\pi_{\phi},\pi_{\text{ref}}\leftarrow\pi_{\theta_{\text{init}}}$
@@
9:  Compute advantages AA with Eq. 5
@@
(2) Unlike common PRMs that produce only step-level rewards, the Implicit PRM provides more fine-grained token-level rewards at no additional cost. This addresses the ambiguity in identifying steps in LLM responses while not introducing extra overhead, making it easy to combine with any RL algorithms for advantage estimation. More discussions on Implicit PRMs are in § C.
@@ 3.2 Advantage Estimation and Policy Update
$A^{i}=r_{o}(\mathbf{y}^{i})-\frac{1}{K-1}\sum_{j\neq i}r_{o}(\mathbf{y}^{j})$ (4)
where $r_{o}(\mathbf{y}^{i})$ denotes the reward of $i$-th response, $K$ is the number of samples for one prompt. The leave-one-out (LOO) baseline helps reduce variances.
@@
Updating policy with PPO loss. We adopt PPO clip surrogate loss for more stable policy updates:
@@
where $\epsilon$ is a clipping parameter. The loss prevents the updated policy from deviating too far from the original distribution, which is the prerequisite of importance sampling.
@@ 3.3 Other Techniques
Initializing PRM with SFT/base model. In practice, we find that the starting policy model itself serves as a decent initialization of PRM, bypassing the PRM training stage. This solves C3 in §2.2 and outperforms a dedicatedly trained PRM, as shown in § 5.1.
@@ 4 Experiments
We first perform supervised finetuning on the base model to get a starter model for RL. Please refer to § D for more details of this stage. All experiments are conducted on $\timesA8$00 GPUs.
@@ 4.1 RL Settings
Rule-based Outcome Verifier. Consistent with recent research that adopts an exact match with ground truth as unhackable rewards (Gao et al., 2024; Lambert et al., 2024; DeepSeek-AI et al., 2025), we define the rule-based ground truth outcome verifiers (OV) for math and coding as follows:
$r_{o}^{\text{math}}(\mathbf{y})$ $=\begin{cases}1,&\text{matched}\\ 0,&\text{otherwise}\end{cases}\quad r_{o}^{\text{code}}(\mathbf{y})=\frac{\sum\text{\#passes}}{\sum\text{\#test cases}}$
@@
Table 1: Detailed results of PRIME and RLOO w/ outcome verifier (OV). At the same 240 steps, the model trained by PRIME is generally better than the model trained by outcome rewards. We also reported avg@16 results in Table 12.
@@
Figure 3: The effect of dense reward. We compare PRIME and RLOO with outcome verifier (OV). PRIME leads to $2.5\times$ sample efficiency (wall clock as X axis can be found in Figure 17) and $6.9\%$ performance improvement. PRIME also substantially outperforms RLOO on downstream tasks.
@@ 4.2 Main Results
As shown in Figure 12 and Table 1, Eurus-2-7B-PRIME achieves substantial improvements on key reasoning benchmarks over the SFT version of the model, leading to 15.1% improvement on average, and over 20% on AMC and AIME competitions. Besides, Eurus-2-7B-PRIME achieves 26.7% pass@1 on AIME 2024, surpassing GPT-4o, Llama-3.1-70B-Instruct, and Qwen2.5-Math-7B-Instruct, demonstrating its excellent reasoning ability. Additional results are in § B.
@@
Performance. We first validate the effect of PRIME with dense rewards compared to RLOO with outcome rewards only. We train this model for 240 steps. For PRIME, we use the same setting and train the model for 592 steps. We plot the training rewards measured by the outcome verifier and test accuracy in Figure 3. Compared with sparse reward, PRIME improves the final rewards by 6.9%, with lower variances. On downstream tasks, PRIME also consistently outperforms OV only setup. Detailed results are listed in Table 1.
Table 2: Step-wise time cost of PRIME and RLOO.
Time(s) Rollout Policy update PRM update Others Sum
PRIME 281.7 156.6 150.9 91.1 680.3
RLOO 282.4 157.9 0 90.4 530.7
[…]
@@ 5 Analysis
Figure 4: Comparison of different PRMs. Online PRM initialized from SFT model achieved the best results. However, using PRMs trained on extra rollouts hurts the performance.
@@ 5.1 Design Choices for the Implicit PRM
Figure 5: Impact of PRM online update. Offline PRM is gradually been overoptimized while online PRMs achieve higher accuracy during training.
SFT model initializes a good PRM. Conventionally, we need to collect data to train RMs and PRMs, and then we can use them in RL. However, the Implicit PRM is a language model, so we can initialize it from any language model with the same tokenizer as the policy model. To investigate whether it is still necessary to train a PRM in advance, we conduct experiments with different PRM initialization strategies: with the SFT model itself and with a specially trained PRM. For the later one, we train EurusPRM from Eurus-2-7B-SFT with additional 500K data generated by Llama3.1 and Qwen2.5 series (data details in § E.5).
@@
5.2 Scaling PRIME with More Compute
(a) PRIME training up to 800 steps.
(b) PRIME training with 16 rollouts per prompt.
Figure 6: RL training with more training steps (Left) and larger rollout numbers (Right).
To validate the scalability of PRIME with increased computational resources, we first conduct an extended training process. Specifically, we conduct RL training for 800 rollout steps (3200 gradient steps) with PRIME and RLOO with outcome-reward only. The training results, shown in Figure 6(a) (Left), reveal that throughout the training, PRIME consistently exhibits stable growth and outperforms the baseline with an improvement of  3.7%. Moreover, we increase the number of responses sampled for each prompt from 4 to 16. The results in Figure 6(a) (Right) show that PRIME achieves non-trivial improvement of approximately 4.4% compared to RLOO.
5.3 PRIME with Other RL Algorithms
As we stated before, PRIME is equally applicable to other RL algorithms beyond RLOO. In this section, we implement PRIME with REINFORCE (Williams, 1992), GRPO (Shao et al., 2024), and PPO (Schulman et al., 2017). Similarly to RLOO, we only modify the advantage estimation functions and leave the clip surrogate loss unchanged. Detailed functions can be found in § B.3. From Figure 8 and Table 4, We show that PRIME boosts these algorithms on both efficiency and performance as it does with RLOO. PRIME contributes consistently regardless of the policy update method, making it a generic algorithm. It indicates that PRIME is a general plug-in for almost any RL algorithm for LLM., which largely extends the use cases of PRIME.
Figure 7: PRIME also generally benefits REINFORCE, GRPO, and PPO.
Figure 8: Comparison of value models and process reward models.
5.4 Value or Reward, How to Use the Implicit PRM?
(1) REINFORCE: $A_{t}=r_{o}(\mathbf{y})$. (2) On top of (1), using a linear-head value model $V$ to calculate the baseline: $A_{t}=r_{o}(\mathbf{y})-V(\mathbf{y}_{<t})$. This is the original PPO in Figure 8 as we set $\gamma=1$ and $\lambda=1$. (3) On top of (1), using values from the Implicit PRM to serve as the baseline: $A_{t}=r_{o}(\mathbf{y})-v_{\phi}(\mathbf{y}_{<t})$. This is equivalent to PPO with its value model being replaced by values from the Implicit PRM when $\gamma=1$ and $\lambda=1$. (4) On top of (1), using process rewards from the Implicit PRM to calculate the return: $A_{t}=r_{o}(\mathbf{y})+\sum_{s=t}^{T}r_{\phi}(y_{s})$. This is exactly the REINFORCE w/ PRIME in Figure 8.
Figure 8 reports the results. Comparing PPO and REINFORCE, we find that an additional value model does not benefit policy performance. Notably, using rewards from the Implicit PRM to calculate returns, which is the default setting in PRIME, greatly outperforms all three baselines. This indicates that PRMs work better than value models in RL for LLMs. both two kinds of value models (PPO value model and Implicit PRM as value model) fall behind reward models.
6 Related Work
RL for LLM Reasoning. In the area of LLMs, reinforcement learning has been widely used for aligning human preferences (Christiano et al., 2017; Ouyang et al., 2022; Cui et al., 2024), but the open-source community mostly adopt imitation learning methods (Yuan et al., 2024a; Yue et al., 2024; Wei et al., 2024; Liu et al., 2024) to enhance the reasoning capabilities of LLMs. Over the past few months, the paradigm gradually shifted. OpenAI o1 (Jaech et al., 2024) first showed the tremendous potential of large-scale RL for reasoning LLMs, and recent works have verified the scaling effect of RL with outcome rewards (DeepSeek-AI et al., 2025; Team et al., 2025). Meanwhile, the role of dense rewards in RL remains underexplored, serving the main focus of PRIME.
Implicit Rewards. Implicit rewards are broadly adopted in LLM alignment (Rafailov et al., 2023; Chen et al., 2024b; Azar et al., 2024; Ethayarajh et al., 2024; Rosset et al., 2024; Chen et al., 2024a).  Rafailov et al. (2024) first showed that optimizing DPO objective learns a Q function implicitly. Zhou et al. (2024) utilized implicit rewards in PPO, and showed the effectiveness of dense implicit rewards.  Yuan et al. (2024b) further extended the conclusion to any loss function optimizing Eq. 3.
7 Conclusion
As the fuel of LLMs, data, will be depleted in the near future, we are entering a new era of experience, which is exemplified by RL (Sutton, 2019). This work develops PRIME, which produces and leverages dense rewards in online RL for LLM reasoning. Throughout the experiments, we validate that PRIME (1) greatly benefits sample efficiency and policy performance, (2) is easy to use with minimum cost, and (3) is a general method that works with broad RL algorithms together.
Ethics Statement
This paper presents PRIME whose goal is to advance the field of reinforcement learning for LLMs. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Reproducibility Statement
We have provided sufficient details to for reproduction, including algorithm pseudocode in Algorithm 1, experiment configurations and hyperparameters in Section 4 and Appendix. We have upload our code in Supplementary Material.
Appendix A Limitations
Due to resource constraints, we only conducted experiments on models up to 32B. Besides the main experiments of PRIME, we ran fewer steps for other ablation experiments, while we conduct comparison under the same step number for fairness.
Appendix B Additional Results
B.1 Reference Model Choice is Flexible